# Polynomial-Time Near-Optimal Estimation over Certain Type-2 Convex Bodies

### v1

@@ 1 Introduction
Many statistical estimation problems are naturally formulated as constrained mean estimation or constrained regression problems. Given observations corrupted by stochastic noise, one seeks an estimator whose target signal is restricted to lie in a known constraint set $K\subset\mathbb{R}^{n}$. Such shape, geometric, or structural constraints arise ubiquitously in nonparametric statistics, signal processing, and machine learning—including monotonicity, convexity, sparsity, total variation, and various Lipschitz or Hölder smoothness classes (e.g., robertson1988order; groeneboom2014nonparametric; guntuboyina2018nonparametric; tsybakov2009introduction).
From a decision-theoretic perspective, a fundamental goal is to characterize the minimax risk over the constraint class $K$. Recent work has made significant progress on this front. In particular, neykov2022minimax derived sharp minimax rates for constrained mean estimation in the Gaussian sequence model, showing that these rates hold for essentially any convex constraint, and depend on the geometric structure of $K$. Furthermore prasadan2024information, relaxed the convexity assumption to a star-shaped assumption which encompasses even more constraints. In the regression setting, prasadan2025characterizingminimaxratenonparametric obtained the corresponding minimax rates under analogous assumptions.
@@ 1.1 Related Work
We now review several related works. In the Gaussian sequence model setting, when the constraint set is an ellipsoid a classical result due to Pinsker pinsker1980optimal shows that linear estimators are optimal. One can implement Pinsker’s estimator in polynomial-time. A similar result holds for linear regression over ellipsoidal constraints goldenshluger2001adaptive; goldenshluger2003optimal. Since ellipsoids are type-2 sets, our framework provides alternative polynomial-time near minimax rate optimal estimators. Notably, there exist ellipsoids over which the least squares estimator is grossly sub-optimal (see zhang2013nearly e.g.). This fact highlights the need to develop computationally efficient algorithms beyond least squares.
zhang2013nearly proposes a near minimax optimal estimator for a linear regression problem with an $\ell_{q}$ constraints on the signal for $q\leq 1$. The author’s algorithm is implementable only when $q=1$. Our work is significantly different, as we require the constraint set to be of type-2 (which the $\ell_{1}$ ball is not). Despite this major difference, there are also some similarities: both our work and zhang2013nearly use approximations of the Kolmogorov widths of the constraint set, as well as Euclidean projections. javanmard2012minimax argue that for bounded symmetric convex polytopes, $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{\infty}\leq 1\}$, defined by some matrix $A\in\mathbb{R}^{m\times n}$, the truncated series estimator [^3] A truncated series estimator uses the optimal Euclidean projection from a carefully selected Kolmogorov width, to project the data on a lower dimensional subspace. is nearly optimal for the Gaussian sequence model problem, provided that $m$ is not too large. However, the authors mention that they do not know how to efficiently implement their estimator. They also formulate a class of constraints of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p>0$, although they only show the near optimality result for the truncated series when $p=\infty$. javanmard2012minimax do conjecture however (see Conjecture 7.1 therein), that the truncated series estimator remains near optimal (within $\log m$ of optimal) for sets of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p\geq 2$. While we cannot verify their conjecture, we do show that our estimator provably works near optimally for any well-balanced constraint of the type $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for any $p\geq 2$ (we lose logarithmic factors in $m$).
Finally, we note a recent related work by kur2022efficient, which develops computationally efficient estimators for multidimensional convex and Lipschitz nonparametric regression. The estimators proposed there are fundamentally different from those studied in this paper, as they are specifically tailored to the structure of convex and Lipschitz constraints, whereas our approach is designed to address constrained estimation problems in a considerably more general setting.
@@ 1.2 Definitions and Notation
The local entropy plays a fundamental role in what follows as neykov2022minimax; prasadan2024information show it determines the minimax rate for the constraint mean estimation problem.
@@ 2 Background and Conditional Algorithm
When $\xi\sim N(0,\sigma^{2}I_{n})$, which is part of our model class, neykov2022minimax and later prasadan2024information showed that when $K$ is a star-shaped set the minimax rate is given by $\eta^{\star 2}\wedge d^{2}$ [^4] In fact, it is easy to see that $\eta^{\star}\lesssim d$ and hence the minimax rate can be written as simply $\eta^{\star 2}$., where $d=\mathrm{diam}(K)$ and $\eta^{\star}:=\eta^{\star}(K)=\sup\{\eta\geq 0:\eta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\eta)\}$ with $M^{\operatorname{loc}}_{K}(\eta)$ being the local entropy of the set $K$. The goal of the present paper is to construct a polynomial-time algorithm which nearly achieves the minimax rate over a range of convex bodies $K$. Observe that we define the minimax rate with $\sigma$ rather than $\underline{\sigma}$ (in the Gaussian case which gives the lower bound, we have $\underline{\sigma}\equiv\sigma$). We will later see (see Section 4) that this can be relaxed to using $\underline{\sigma}$ in the definition of $\eta^{\star}$, but it requires having more than a single sample.
@@
In other words, $(\mathbb{R}^{n},\rho_{K})$ is a Banach space of type-2. Note that from the above definition it is clear that $T_{2}(K)\geq 1$ always. We will use this fact in several places. Going back to the “small” type-2 constant — here by small we mean that $T_{2}(X)$ is a slowly growing function of the parameters ($n,d,\sigma$). For instance $T_{2}(K)=\sqrt{\log n}$ is a small constant (in contrast to $T_{2}(K)=\sqrt{n}$ which is a large constant). Our third assumption is that we have an oracle for computing $\rho_{K}$ (which implies we have a (weak) membership oracle for $K$; see grotschel2012geometric for a definition and properties of (weak) membership oracles). Our fourth assumption is that there are known constants $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$. We note here that the complexity of our algorithms depends polynomially on $\log 1/r$ and $\log R$, so in principle $r,R$ can be exponentially small/large with $n$, and hence we do not view this as a severe restriction. We will refer to such a set $K$ (i.e., a set sandwiched between two Euclidean balls) as well-balanced. Also, observe that $R=d/2$ always works, provided that $d$ is known, since the set $K$ is origin symmetric. Otherwise we must have $R\geq d/2$.
Since the space $(\mathbb{R}^{n},\|\cdot\|_{2})$ is self-dual with $T_{2}(B_{2})$ constant equal to $1$, by a classical result of B. Carl, (carl1985inequalities, see Theorem 5), we have up to an absolute constant factor:
@@
Observe that now that for any $\delta\geq 2\eta^{*}$ we have $\log M^{\operatorname{loc}}(\delta)\leq\log M^{\operatorname{loc}}(2\eta^{\star})\leq 4\eta^{\star 2}/\sigma^{2}$ (see neykov2022minimax Lemma II.8 where it is shown that for convex sets the map $\delta\mapsto\log M^{\operatorname{loc}}(\delta)$ is non-increasing). Set $\kappa^{-1/2}\cdot\tilde{d}/T_{2}(K)=\Upsilon$. Under our assumptions we have that $\Upsilon\geq 2\eta^{*}$. Next, by Lemma 3 of yang1999information we know that
@@
Unfortunately, we do not know the optimal projection that yields a small Kolmogorov width. In addition, even calculating the Kolmogorov width $d_{m}(K)$ precisely in general is likely NP hard – see, for instance, brieden2002geometric where the author shows that calculating $d_{0}(K)$ for (even symmetric) convex polytopes cannot be computed in polynomial-time within a factor of $1.090$, unless P=NP. Thus, the fact (2.3) appears useless at first glance. However, since we are interested in having a minimax optimal estimator up to “small” factors, we do not need to find the precise optimal Kolmogorov projection. Taking inspiration from varadarajan2007approximating, we propose to solve the following semi-definite program (SDP) relaxation of the Kolmogorov width problem
@@
As observed in varadarajan2007approximating, the optimal value $X^{\star}$ that this SDP returns, satisfies $p^{\top}X^{\star}p\leq d^{2}_{m}(K)$ for all $p\in K$. The proof of this fact is elementary, and we include a proof below for convenience of the reader. The proof also helps motivate program (2.6).
@@
This is equivalent to $p^{\top}P_{n-m}p\leq d^{2}_{m}(K)$. Note that $P_{n-m}$ is a feasible point to the SDP hence the result follows. ∎
Notably, solving the SDP requires maximizing a quadratic form over a convex body (cf. (2.6)), a formidable task on its own. The authors of varadarajan2007approximating do not encounter this issue, since their body $K$ is a convex polytope with finitely many vertices, making the quadratic maximization tractable (the maximum is attained at a vertex and can be found by enumerating all vertices). In our setting, no such guarantee holds. We will now assume that there exists an oracle $\mathcal{O}_{K}$ which takes a symmetric psd matrix $X$ and outputs a vector $\mathcal{O}_{K}(X)\in K$ such that
@@ 2 Projection onto the capped simplex $S_{k}=\{w:0\leq w_{i}\leq 1,\sum_{i}w_{i}=k\}$
where we used (2.8) and the fact that $\langle g^{(k)},X^{\star}\rangle\leq f(X^{\star})$. We will now repeat the original proof of the sub-gradient descent due to Shor shor2012minimization; boyd2004convex, using the near sub-gradient property above. Since Euclidean projection on a convex set shrinks the distance to any point in the set, we have
@@
$f(X^{(\text{best})})$ $\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+f(X^{\star}))\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+d_{m}(K)^{2})\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+\kappa^{-1}\cdot\tilde{d}^{2}/c^{2})$
@@
Suppose $\tilde{d}\geq C\sigma$. Then the estimator $A^{\star\star}Y$ of $A^{\star\star}\mu$ satisfies
@@
Applying the Hanson-Wright inequality rudelson2013hanson, we obtain:
@@
The fact that $\tilde{d}\gtrsim\sigma$ implies that $m\asymp\tilde{d}^{2}/(C^{2}\sigma^{2})$, where $\asymp$ means equality up to absolute constant factors (i.e., $\gtrsim$ and $\lesssim$ both hold). Now $\|A^{\star\star 2}\|_{\operatorname{op}}\leq 1$. In addition $\|A^{\star\star 2}\|^{2}_{F}=\operatorname{tr}((I-X^{\star\star})^{2})\leq\operatorname{tr}(I-X^{\star\star})=m$.
Thus we can select $t=\tilde{d}^{2}/c^{2}\gtrsim m\sigma^{2}$ to claim that (2.9) holds with high probability, i.e., probability at least $1-\exp(-c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2}))$. Comparing $\tilde{d}^{4}/(\sigma^{4}m)\lesssim\tilde{d}^{2}/\sigma^{2}$ yields $\tilde{d}^{2}\lesssim\sigma^{2}m$ which implies $c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2})\asymp\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2})$ for some $C^{\prime}>0$. So that the bound is always $1-\exp(-\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2}))$. Since the mean of $\|A^{\star\star}\xi\|_{2}^{2}$ is $\underline{\sigma}^{2}\operatorname{tr}(A^{\star\star 2})=\underline{\sigma}^{2}m\lesssim\sigma^{2}\tilde{d}^{2}/(C^{2}\sigma^{2})$ we complete the proof. ∎
@@
Lemma 2.7 is the reason why we assumed that the noise $\xi$ has independent components, so that we can apply the Hanson-Wright inequality. We note that the formula for the minimax rate remains unchanged even if the noise has dependent entries prasadan2024information. In Section 4 we will see that if we have more samples the noise can be arbitrary provided that its covariance matrix has a bounded operator norm.
@@
Suppose we have an estimator of $\mu$ such that $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ happens with high probability, for some $\tilde{d}_{j}\in\mathbb{R}_{+}$ such that $\tilde{d}_{j}\geq 2r\vee C\sigma$. Let $K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Obtain $m$ and the matrix $A_{j}^{\star\star}$ from Lemma 2.1 and Algorithm 2 for the set $K_{(j)}$ [^6] We explain later that the oracle $\mathcal{O}_{K_{(j)}}$ can also be run in polynomial-time, provided that the oracle $\mathcal{O}_{K}$ can be run in polynomial-time., assuming that the conditions of Lemma 2.1 hold. Then the estimator $\tilde{\mu}_{j+1}=(A^{\star\star}_{j}Y-A^{\star\star}\hat{\mu}_{j})/2$ of $(\mu-\hat{\mu}_{j})/2$ used in Lemma 2.7) satisfies:
@@
On the event $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ we have that, $\nu=(\mu-\hat{\mu}_{j})/2\in K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Recall that $A_{j}^{\star\star}=I-X_{j}^{\star\star}$, and let $X_{j}^{\star\star}=V\Lambda V^{\top}$ be the eigendecomposition of $X_{j}^{\star\star}$.
@@
$=\sum_{i}\frac{\lambda_{i}^{2}}{(1+\sqrt{1-\lambda_{i}})^{2}}(v_{i}^{\top}\nu)^{2}\leq\sum_{i}\lambda_{i}(v_{i}^{\top}\nu)^{2}=\nu^{\top}X_{j}^{\star\star}\nu\leq(\tilde{d}_{j}/c)^{2},$
where we used $0\leq\lambda_{i}\leq 1$. Next, $\|\tilde{\mu}_{j+1}-(\mu-\hat{\mu}_{j})/2\|_{2}\leq\|A_{j}^{\star\star}Y/2-A_{j}^{\star\star}\mu/2\|_{2}+\|A_{j}^{\star\star}(\mu-\hat{\mu}_{j})/2-(\mu-\hat{\mu}_{j})/2\|_{2}\leq\tilde{d}_{j}/(2T)+(\tilde{d}_{j}/c)\leq\tilde{d}_{j}/L$, with probability at least $1-\exp(-\tilde{d}_{j}^{2}/(C^{{}^{\prime}2}\sigma^{2}))-(1+\tilde{d}^{3}/\sigma^{3})$ by Lemma 2.7. This completes the proof. ∎

### v2

@@ 1 Introduction
Many statistical estimation problems are naturally formulated as constrained mean estimation or constrained regression problems. Given observations corrupted by stochastic noise, one seeks an estimator whose target signal is restricted to lie in a known constraint set $K\subset\mathbb{R}^{n}$. Such shape, geometric, or structural constraints arise ubiquitously in nonparametric statistics, signal processing, and machine learning—including monotonicity, convexity, sparsity, total variation, and various Lipschitz or Hölder smoothness classes (e.g., Robertson (1988); Groeneboom and Jongbloed (2014); Guntuboyina and Sen (2018); Tsybakov (2009)).
From a decision-theoretic perspective, a fundamental goal is to characterize the minimax risk over the constraint class $K$. Recent work has made significant progress on this front. In particular, Neykov (2022) derived sharp minimax rates for constrained mean estimation in the Gaussian sequence model, showing that these rates hold for essentially any convex constraint, and depend on the geometric structure of $K$. Furthermore Prasadan and Neykov (2024), relaxed the convexity assumption to a star-shaped assumption which encompasses even more constraints. In the regression setting, Prasadan and Neykov (2025a) obtained the corresponding minimax rates under analogous assumptions.
@@ 1.1 Related Work
We now review several related works. In the Gaussian sequence model setting, when the constraint set is an ellipsoid a classical result due to Pinsker Pinsker (1980) shows that linear estimators are optimal. One can implement Pinsker’s estimator in polynomial-time. A similar result holds for linear regression over ellipsoidal constraints Goldenshluger and Tsybakov (2001, 2003). Since ellipsoids are type-2 sets, our framework provides alternative polynomial-time near minimax rate optimal estimators. Notably, there exist ellipsoids over which the least squares estimator is grossly sub-optimal (see Zhang (2013) e.g.). This fact highlights the need to develop computationally efficient algorithms beyond least squares.
Zhang (2013) proposes a near minimax optimal estimator for a linear regression problem with an $\ell_{q}$ constraints on the signal for $q\leq 1$. The author’s algorithm is implementable only when $q=1$. Our work is significantly different, as we require the constraint set to be of type-2 (which the $\ell_{1}$ ball is not). Despite this major difference, there are also some similarities: both our work and Zhang (2013) use approximations of the Kolmogorov widths of the constraint set, as well as Euclidean projections. Javanmard and Zhang (2012) argue that for bounded symmetric convex polytopes, $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{\infty}\leq 1\}$, defined by some matrix $A\in\mathbb{R}^{m\times n}$, the truncated series estimator [^3] A truncated series estimator uses the optimal Euclidean projection from a carefully selected Kolmogorov width, to project the data on a lower dimensional subspace. is nearly optimal for the Gaussian sequence model problem, provided that $m$ is not too large. However, the authors mention that they do not know how to efficiently implement their estimator. They also formulate a class of constraints of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p>0$, although they only show the near optimality result for the truncated series when $p=\infty$. Javanmard and Zhang (2012) do conjecture however (see Conjecture 7.1 therein), that the truncated series estimator remains near optimal (within $\log m$ of optimal) for sets of the type $\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for $p\geq 2$. While we cannot verify their conjecture, we do show that our estimator provably works near optimally for any well-balanced constraint of the type $K=\{x\in\mathbb{R}^{n}:\|Ax\|_{p}\leq 1\}$ for any $p\geq 2$ (we lose logarithmic factors in $m$).
Finally, we note a recent related work by Kur and Putterman (2022), which develops computationally efficient estimators for multidimensional convex and Lipschitz nonparametric regression. The estimators proposed there are fundamentally different from those studied in this paper, as they are specifically tailored to the structure of convex and Lipschitz constraints, whereas our approach is designed to address constrained estimation problems in a considerably more general setting.
@@ 1.2 Definitions and Notation
The local entropy plays a fundamental role in what follows as Neykov (2022); Prasadan and Neykov (2024) show it determines the minimax rate for the constraint mean estimation problem.
@@ 2 Background and Conditional Algorithm
When $\xi\sim N(0,\sigma^{2}I_{n})$, which is part of our model class, Neykov (2022) and later Prasadan and Neykov (2024) showed that when $K$ is a star-shaped set the minimax rate is given by $(\eta^{\star 2}\wedge d^{2})$ [^4] In fact, it is easy to see that $\eta^{\star}\lesssim d$ and hence the minimax rate can be written as simply $\eta^{\star 2}$., where $d=\mathrm{diam}(K)$ and $\eta^{\star}:=\eta^{\star}(K)=\sup\{\eta\geq 0:\eta^{2}/\sigma^{2}\leq\log M^{\operatorname{loc}}_{K}(\eta)\}$ with $M^{\operatorname{loc}}_{K}(\eta)$ being the local entropy of the set $K$. The goal of the present paper is to construct a polynomial-time algorithm which nearly achieves the minimax rate over a range of convex bodies $K$. Observe that we define the minimax rate with $\sigma$ rather than $\underline{\sigma}$ (in the Gaussian case which gives the lower bound, we have $\underline{\sigma}\equiv\sigma$). We will later see (see Section 4) that this can be relaxed to using $\underline{\sigma}$ in the definition of $\eta^{\star}$, but it requires having more than a single sample.
@@
In other words, $(\mathbb{R}^{n},\rho_{K})$ is a Banach space of type-2. Note that from the above definition it is clear that $T_{2}(K)\geq 1$ always. We will use this fact in several places. Going back to the “small” type-2 constant — here by small we mean that $T_{2}(X)$ is a slowly growing function of the parameters ($n,d,\sigma$). For instance $T_{2}(K)=\sqrt{\log n}$ is a small constant (in contrast to $T_{2}(K)=\sqrt{n}$ which is a large constant). Our third assumption is that we have an oracle for computing $\rho_{K}$ (which implies we have a (weak) membership oracle for $K$; see Grötschel et al. (2012) for a definition and properties of (weak) membership oracles). Our fourth assumption is that there are known constants $r,R\in\mathbb{R}_{+}$ such that $rB_{2}\subseteq K\subseteq RB_{2}$. We note here that the complexity of our algorithms depends polynomially on $\log 1/r$ and $\log R$, so in principle $r,R$ can be exponentially small/large with $n$, and hence we do not view this as a severe restriction. We will refer to such a set $K$ (i.e., a set sandwiched between two Euclidean balls) as well-balanced. Also, observe that $R=d/2$ always works, provided that $d$ is known, since the set $K$ is origin symmetric. Otherwise we must have $R\geq d/2$.
Since the space $(\mathbb{R}^{n},\|\cdot\|_{2})$ is self-dual with $T_{2}(B_{2})$ constant equal to $1$, by a classical result of B. Carl, (Carl, 1985, see Theorem 5), we have up to an absolute constant factor:
@@
Observe now that for any $\delta\geq 2\eta^{*}$ we have $\log M^{\operatorname{loc}}(\delta)\leq\log M^{\operatorname{loc}}(2\eta^{\star})\leq 4\eta^{\star 2}/\sigma^{2}$ (see Neykov (2022) Lemma II.8 where it is shown that for convex sets the map $\delta\mapsto\log M^{\operatorname{loc}}(\delta)$ is non-increasing). Set $\kappa^{-1/2}\cdot\tilde{d}/T_{2}(K)=\Upsilon$. Under our assumptions we have that $\Upsilon\geq 2\eta^{*}$. Next, by Lemma 3 of Yang and Barron (1999) we know that
@@
Unfortunately, we do not know the optimal projection that yields a small Kolmogorov width. In addition, even calculating the Kolmogorov width $d_{m}(K)$ precisely in general is likely NP hard – see, for instance, Brieden (2002) where the author shows that calculating $d_{0}(K)$ for (even symmetric) convex polytopes cannot be computed in polynomial-time within a factor of $1.090$, unless P=NP. Thus, the fact (2.3) appears useless at first glance. However, since we are interested in having a minimax optimal estimator up to “small” factors, we do not need to find the precise optimal Kolmogorov projection. Taking inspiration from Varadarajan et al. (2007), we propose to solve the following semi-definite program (SDP) relaxation of the Kolmogorov width problem
@@
As observed in Varadarajan et al. (2007), the optimal value $X^{\star}$ that this SDP returns, satisfies $p^{\top}X^{\star}p\leq d^{2}_{m}(K)$ for all $p\in K$. The proof of this fact is elementary, and we include a proof below for convenience of the reader. The proof also helps motivate program (2.6).
@@
This is equivalent to $\sup_{p\in K}p^{\top}P_{n-m}p\leq d^{2}_{m}(K)$. Note that $P_{n-m}$ is a feasible point to the SDP hence the result follows. ∎
Notably, solving the SDP requires maximizing a quadratic form over a convex body (cf. (2.6)), a formidable task on its own. The authors of Varadarajan et al. (2007) do not encounter this issue, since their body $K$ is a convex polytope with finitely many vertices, making the quadratic maximization tractable (the maximum is attained at a vertex and can be found by enumerating all vertices). In our setting, no such guarantee holds. We will now assume that there exists an oracle $\mathcal{O}_{K}$ which takes a symmetric psd matrix $X$ and outputs a vector $\mathcal{O}_{K}(X)\in K$ such that
@@ 2 Projection onto the capped simplex $S_{k}=\{w:0\leq w_{i}\leq 1,\sum_{i}w_{i}=k\}$
where we used (2.8) and the fact that $\langle g^{(k)},X^{\star}\rangle\leq f(X^{\star})$. We will now repeat the original proof of the sub-gradient descent due to Shor Shor (2012); Boyd and Vandenberghe (2004), using the near sub-gradient property above. Since Euclidean projection on a convex set shrinks the distance to any point in the set, we have
@@
$f(X^{(\text{best})})$ $\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+f(X^{\star}))\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+d_{m}(K)^{2})\leq\kappa\cdot(\kappa^{-1}\cdot\tilde{d}^{2}/C^{2}+L\cdot\kappa^{-1}\cdot\tilde{d}^{2}/c^{2})$
@@
Supopse $\tilde{d}\gtrsim\sigma$. Then, the estimator $A^{\star\star}Y$ of $A^{\star\star}\mu$ satisfies
@@
Applying the Hanson-Wright inequality Rudelson and Vershynin (2013), we obtain:
@@
Note that $\|A^{\star\star 2}\|_{\operatorname{op}}\leq 1$. In addition $\|A^{\star\star 2}\|^{2}_{F}=\operatorname{tr}((I-X^{\star\star})^{2})\leq\operatorname{tr}(I-X^{\star\star})=m$.
Hence we can select $t=\tilde{d}^{2}/c^{2}\gtrsim m\sigma^{2}$ to claim that (2.9) holds with high probability, i.e., probability at least $1-\exp(-c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2}))$. Comparing $\tilde{d}^{4}/(\sigma^{4}m)\gtrsim\tilde{d}^{2}/\sigma^{2}$ yields $\tilde{d}^{2}\gtrsim\sigma^{2}m$ which is true as $m=\lceil\tilde{d}^{2}/(C^{2}\sigma^{2})\rceil\wedge n$ and $\tilde{d}\gtrsim\sigma$. This implies $c^{\prime}\tilde{d}^{4}/(\sigma^{4}m)\wedge\tilde{d}^{2}/(c^{2}\sigma^{2})\asymp\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2})$ for some $C^{\prime}>0$. So that the bound is always $1-\exp(-\tilde{d}^{2}/(C^{{}^{\prime}2}\sigma^{2}))$. Since the mean of $\|A^{\star\star}\xi\|_{2}^{2}$ is $\underline{\sigma}^{2}\operatorname{tr}(A^{\star\star 2})=\underline{\sigma}^{2}m\lesssim\sigma^{2}\tilde{d}^{2}/(C^{2}\sigma^{2})$ we complete the proof. ∎
@@
Lemma 2.7 is the reason why we assumed that the noise $\xi$ has independent components, so that we can apply the Hanson-Wright inequality. We note that the formula for the minimax rate remains unchanged even if the noise has dependent entries Prasadan and Neykov (2024). In Section 4 we will see that if we have more samples the noise can be arbitrary provided that its covariance matrix has a bounded operator norm.
@@
Suppose we have an estimator of $\mu$ such that $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ happens with high probability, for some $\tilde{d}_{j}\in\mathbb{R}_{+}$ such that $\tilde{d}_{j}\geq 2r\vee C\sigma$. Let $K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Obtain $m$ and the matrix $A_{j}^{\star\star}$ from Lemma 2.1 and Algorithm 2 for the set $K_{(j)}$ [^6] We explain later that the oracle $\mathcal{O}_{K_{(j)}}$ can also be run in polynomial-time, provided that the oracle $\mathcal{O}_{K}$ can be run in polynomial-time., assuming that the conditions of Lemma 2.1 hold. Then the estimator $\tilde{\mu}_{j+1}=(A^{\star\star}_{j}Y-A_{j}^{\star\star}\hat{\mu}_{j})/2$ of $(\mu-\hat{\mu}_{j})/2$ used in Lemma 2.7 satisfies:
@@
On the event $\|\hat{\mu}_{j}-\mu\|_{2}\leq\tilde{d}_{j}$ we have that, $\nu=(\mu-\hat{\mu}_{j})/2\in K_{(j)}=K\cap B(0,\tilde{d}_{j}/2)$. Recall that $A_{j}^{\star\star}=(I-X_{j}^{\star\star})^{1/2}$, and let $X_{j}^{\star\star}=V\Lambda V^{\top}$ be the eigendecomposition of $X_{j}^{\star\star}$.
@@
$=\sum_{i}\frac{\lambda_{i}^{2}}{(1+\sqrt{1-\lambda_{i}})^{2}}(v_{i}^{\top}\nu)^{2}\leq\sum_{i}\lambda_{i}(v_{i}^{\top}\nu)^{2}=\nu^{\top}X_{j}^{\star\star}\nu\lesssim(\tilde{d}_{j}/c)^{2},$
where we used $0\leq\lambda_{i}\leq 1$. Next, $\|\tilde{\mu}_{j+1}-(\mu-\hat{\mu}_{j})/2\|_{2}\leq\|A_{j}^{\star\star}Y/2-A_{j}^{\star\star}\mu/2\|_{2}+\|A_{j}^{\star\star}(\mu-\hat{\mu}_{j})/2-(\mu-\hat{\mu}_{j})/2\|_{2}\lesssim\tilde{d}_{j}/(2T)+(\tilde{d}_{j}/c)\leq\tilde{d}_{j}/L$, with probability at least $1-\exp(-\tilde{d}_{j}^{2}/(C^{{}^{\prime}2}\sigma^{2}))-(1+\tilde{d}^{3}/\sigma^{3})$ by Lemma 2.7. This completes the proof. ∎